### v2

# Dependency-Aware Semi-Structured Sparsity: Declining Roles of Outliers in Pruning GLU-based LLMs

Abstract
The rapid growth in the scale of Large Language Models (LLMs) has led to significant computational and memory costs, making model compression techniques such as network pruning increasingly crucial for their efficient deployment. Recent LLMs such as LLaMA2 and Mistral have adopted GLU-based MLP architectures. However, current LLM pruning strategies are primarily based on insights from older LLM architectures, necessitating a reevaluation of these strategies to suit the new architectural characteristics. Contrary to traditional beliefs, we find that outliers play a diminished role in the input projections of GLU-based MLPs. Leveraging this new insight, we propose Dependency-aware Semi-structured Sparsity (DaSS), a novel pruning method for GLU-based LLMs. DaSS balances the flexibility of unstructured pruning and the structural consistency of dependency-based structured pruning by considering both of weight magnitude and corresponding intermediate activation norms in weight pruning metric. Empirical evaluations on the Mistral, Gemma, and LLaMA2 model families demonstrate the consistent effectiveness of DaSS in the prevailing GLU variants.
Dependency-Aware Semi-Structured Sparsity:
Declining Roles of Outliers in Pruning GLU-based LLMs
Zhiyu Guo, Hidetaka Kamigaito, Taro Watanabe Nara Institute of Science and Technology {guo.zhiyu.fy1, kamigaito.h, taro}@is.naist.jp
1 Introduction
Recent years have witnessed the great success of Large Language Models (LLMs) across various challenging tasks, such as mathematical reasoning, code generation. However, the practical use of these models for inference has faced a major obstacle due to the substantial computational resources they consume. To tackle this, many of the key developments up to now have revolved around weight quantization. It is possible to quantize LLMs down to 4 bits per weight with little impact on accuracy, which aids in memory reduction and speeds up inference Lin et al. (2024). Nonetheless, accuracy recovery becomes problematic when quantizing to around 3 bits per weight with existing methods Dettmers et al. (2024); Egiazarian et al. (2024).
A complementary method is neural network pruning Han et al. (2015b), which can be combined with quantization to further improve the inference efficiency of LLMs Kurtic et al. (2023); Frantar and Alistarh (2023). Pruning can be categorized into two main approaches: unstructured pruning Sun et al. (2024); Frantar and Alistarh (2023), which involves the removal of specific weights, and structured pruning Ma et al. (2023), which entails the removal of complete rows or columns of weights. In contrast to structured pruning, which struggles with performance in LLMs even at low sparsity levels, unstructured pruning methods like SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2024) exhibit promising results without additional retraining, and achieves practical speedup in both CPU and GPU through the recent engineering advancements Agarwalla et al. (2024), they can also benefit in reducing hallucinations of LLMs Chrysostomou et al. (2024).
(a) Dependency-based Structured Sparsity
(b) Wanda Unstructured Sparsity
(c) Dependency-aware Semi-structured Sparsity
Figure 1: Illustration of Dependency-aware Semi-structured Sparsity (DaSS). In (a) dependency-based structured pruning Ma et al. (2023), all the weights connecting to the same intermediate neuron are removed or remain simultaneously. In (b) Wanda unstructured pruning Sun et al. (2024), it assigns greater emphasis to the weights corresponding to large input activations. For Gate-Proj and UP-Proj, the same number of weights are removed for each MLP neuron, regardless of whether some neurons have much larger activation norms. For Down-Proj, the weights corresponding to larger activation norms are more likely to be pruned. This can lead to a structural mismatch. In (c), all the weights corresponding to large intermediate activations are more likely to be reserved.
Modern LLMs, such as LLaMA2 Touvron et al. (2023) and Mistral Jiang et al. (2023), have adopted several architectural changes, including the use of GLU (Gated Linear Unit) variants (e.g., SwiGLU) Shazeer (2020) in the MLP modules, and grouped-query attention (GQA) Ainslie et al. (2023). As the GLU-based MLP module accounts for more than 80% of the parameters in LLMs that use GQA  [^1] In LLaMA2-70B, the dimension of key and value is $\frac{1}{8}d$, MLP intermediate dimension is $\frac{7}{2}d$., its pruning emerges as a pivotal factor in determining the overall compression efficacy of LLMs. Notably, prevailing LLM pruning methods such as SparseGPT and Wanda draw inspiration from Llm.int8() Dettmers et al. (2022), which primarily derived its insights from older LLMs like OPT Zhang et al. (2022). However, with the shift in architectural design towards GLU-based MLP modules, it becomes essential to re-evaluate and potentially redefine our LLM pruning strategies to accommodate these new characteristics. Therefore, a timely question arises: which specific properties of GLU-based MLP demand the development of new LLM pruning approaches?
The prevailing belief has been that outlier features are critically important in LLMs Dettmers et al. (2022), and Wanda is a prime example of an outlier-dependent method. Our initial investigation focuses on a more nuanced examination of the effectiveness of Wanda. Sun et al. (2024) suggests that input-balanced magnitude pruning, which prunes the same amount of weights for each input channel, can achieve performance only slightly worse than Wanda. We divide the GLU-based MLP module into two groups: input projections (Gate-Proj and Up-Proj) and output projection (Down-Proj). We compare the effectiveness of input-balanced magnitude pruning and Wanda on these two groups separately. For the input projections, we surprisingly find that input-balanced magnitude pruning performs better than Wanda when sparsity level is higher than 50%. This suggests that outliers play a diminished role in the input projections of GLU-based MLP, contrary to common assumptions. However, for the output projection, Wanda still performs better.
In dependency-aware structured pruning, it’s critical to consider that pruned parameters have dependencies with other parameters, owing to their interconnected nature Ma et al. (2023); Fang et al. (2023). For GLU-based MLP pruning, all the weights connected to the same intermediate neuron should be preserved or pruned simultaneously.
Leveraging our newfound insight, and drawing inspiration from dependency-based structured pruning method, we introduce a new LLM pruning paradigm, namely, Dependency-aware Semi-structured Sparsity (DaSS). This approach is specifically designed to navigate the middle ground between the flexibility of unstructured pruning and the structural consistency of dependency-based structured pruning. We present a new MLP pruning metric that assesses each weight’s importance based on the product of its magnitude and the norm of the corresponding MLP intermediate activations. Our proposed DaSS method, illustrated in Figure 1(c), embodies a semi-structured pattern that retains a degree of the adaptability inherent in unstructured pruning while incorporating the dependency-aware aspect of structured pruning. This balance allows for more precise pruning.
We perform extensive experiments on LLaMA2  Touvron et al. (2023), Gemma  Team et al. (2024), and Mistral  Jiang et al. (2023) to evaluate DaSS across various tasks from language modeling, 5 commonsense reasoning tasks. In achieving hardware-friendly N:M sparsity patterns, DaSS consistently excels beyond the existing LLM pruning methods SparseGPT and Wanda, while maintaining the computational efficiency akin to Wanda. Moreover, DaSS demonstrates consistent effectiveness in all the prevalent GLU variants, including SwiGLU, GeGLU and ReGLU. Impressively, even without weight updates used in SparseGPT, we observe significant improvements achieved by DaSS over Wanda with LLaMA2-70B in 5 commonsense reasoning tasks, with mean accuracy improvement of 3.63 and 12.24 at sparsity levels of 70% and 80%, respectively. Note that the MLP input projections account for around 50% parameters in GQA-based LLMs, DaSS prunes them without using input activation, challenging the traditional wisdom that outliers play a pivotal role in LLM compression. We hope our fresh insights can also motivate more nuanced GLU-specific LLM quantization strategies.
2 Related Work
Pruning LLM.
Neural network pruning in LLM can be broadly categorized into two groups: structured pruning Ma et al. (2023); Xia et al. (2023a) and unstructured pruning Frantar and Alistarh (2023); Sun et al. (2024). Ma et al. (2023) proposes a dependency detection algorithm to detect and prune non-critical grouped structures. Although structured pruning can have better hardware efficiency, the accuracy drops a lot even at a low compression rate, they typically require re-trianing to recover accuracy. Unstructured pruning can yield a higher compression rate and achieve acceleration on both CPUs and GPUs through the recent software and handware advancement Agarwalla et al. (2024). SparseGPT Frantar and Alistarh (2023) leverages the Hessian inverse for pruning and reduces reconstruction error of dense and sparse weights by subsequent weight updates. Wanda Sun et al. (2024) employs an efficient method that augments input activations into weight magnitudes, and matches the performance of SparseGPT at medium sparsity. OWL Yin et al. (2024) explores non-uniform layer-wise sparsity for pruning LLMs, however, it is less practical to be converted into N:M sparsity. Our work incorporates dependency information into unstructured pruning, achieving a novel pruning paradigm.
Outlier-dependent LLM Compression.
Outlier features, defined as features with magnitudes substantially larger than others, are a notable characteristic of LLMs Dettmers et al. (2022). Despite making up only a small fraction of all feature dimensions, these outliers play a critical role in attention and predictive performance. Such observation has motivated the development of LLM-specific quantization methods Dettmers et al. (2022); Xiao et al. (2023); Lin et al. (2024); Ashkboos et al. (2023) to handle outliers more effectively. Wanda Sun et al. (2024) and OWL Yin et al. (2024) broaden these insights, revealing that outlier features are significant in deciding weight importance when pruning LLMs. Our analysis diverges from conventional wisdom by demonstrating that, in the context of GLU-based MLP input projections, the influence of outlier features is not as pronounced as previously assumed, prompting a reevaluation of their role in LLM compression strategies.
3 Preliminary Exploration
3.1 Wanda Pruning Method
In the context of LLM pruning, we denote a linear layer weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ and input activations $\mathbf{X}\in\mathbb{R}^{L\times d_{\text{in}}}$, where $d_{\text{in}}$, $d_{\text{out}}$ and $L$ represent weight input dimension, weight output dimension, and input sequence length, respectively. Wanda Sun et al. (2024) evaluates weight importance through the multiplication of the weight magnitude with the norm of the input feature. Specifically, the importance score for a given weight $\mathbf{W}_{i,j}$ is determined as follows:
$\mathbf{I}_{i,j}=\left|\mathbf{W}_{i,j}\right|\cdot\left\|\mathbf{X}_{j}\right\|_{2}$ (1)
Here, $\mathbf{I}_{i,j}$ is the weight importance score for a given weight $\mathbf{W}_{i,j}$ in the weight matrix. $\left\|\mathbf{X}_{j}\right\|_{2}$ represents the $\ell_{2}$ norm of the $j^{th}$ feature of input $\mathbf{X}$, which is calculated across all $L$ tokens to produce a scalar. In Wanda pruning, importance scores for weights are compared for each output individually, corresponding to each row in the matrix $\mathbf{W}$. We call this as output-balanced granularity.
3.2 GLU Variants for Transformer
Gated Linear Units (GLU) Dauphin et al. (2017) are formed by the element-wise multiplication of two linear projections, with a sigmoid function applied to one projection before the multiplication. Shazeer (2020) suggests an alternative design for the Transformer’s MLP layer that incorporates GLU variants, effectively replacing the conventional first linear transformation and activation function. More formally, we denote the $d_{\text{hidden}}$ as the dimension of the model’s hidden state and $d_{\text{int}}$ as intermediate dimension of the MLP module. For $\mathbf{x}\in\mathbb{R}^{d_{\text{hidden}}}$, $\mathbf{W_{1}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{int}}}$, $\mathbf{W_{2}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{int}}}$, $\mathbf{W_{3}}\in\mathbb{R}^{d_{\text{int}}\times d_{\text{hidden}}}$, and activation function $\sigma(\cdot)$, each MLP layer produces $\mathbf{z}\in\mathbb{R}^{d_{\text{hidden}}}$ by evaluating
$\mathbf{y}=\sigma(\mathbf{x}\mathbf{W_{1}})\otimes\mathbf{x}\mathbf{W_{2}}$ (2)
$\mathbf{z}=\mathbf{y}\mathbf{W_{3}}$ (3)
We call $\mathbf{W_{1}},\mathbf{W_{2}},\mathbf{W_{3}}$ linear projections as Gate-Proj, Up-Proj, and Down-Proj, respectively. GLU variants that use Swish Ramachandran et al. (2017), GeLU Hendrycks and Gimpel (2016), and ReLU Glorot et al. (2011) activation functions in Eq.(2) are called SwiGLU, GeGLU, and ReGLU, respectively. SwiGLU is most widely used in the recent LLMs.
3.3 Revisiting the Effectiveness of Wanda
Sun et al. (2024) conducted ablation studies among different granularities of magnitude pruning, including input-balanced granularity, which sorts weights connecting to the same input neuron. Compared with Wanda pruning, input-balanced magnitude pruning is just slightly worse.
Here, we further investigate the effectiveness of input-balanced magnitude pruning for the MLP module. We divide the GLU-based MLP module into two groups, input projections (Gate-Proj and Up-Proj) and output projection (Down-Proj). We prune those two groups separately and compare Wanda pruning with input-balanced magnitude pruning. Table 1 shows that input-balanced magnitude pruning is surprisingly better than Wanda pruning at higher sparsity ratios for input projections pruning. For the output projection, input-balanced magnitude pruning is consistently worse than Wanda pruning. This suggests the diminished importance of outlier features within input projections compared to the output projection. However, in the older GeLU-based MLP expansion projection, we observe Wanda can achieve much better performance as shown in Appendix A.3, suggesting this is a special property of GLU variants.
Method Up+Gate Projection Down Projection
50% 60% 70% 50% 60% 70%
Wanda 6.05 7.12 12.42 5.76 6.11 6.97
Magnitude 6.12 7.10 10.52 5.87 6.56 9.03
Table 1: Comparison of Wanda pruning and Input-balanced Magnitude pruning, perplexity with LLaMA2-7B on WikiText
3.4 Output-balanced Pruning Limitations
N:M sparsity pattern offers notable speed improvements on recent NVIDIA GPUs Mishra et al. (2021); Kurtic et al. (2023). It is specified that every group of M consecutive weights must include N zeros. SparseGPT Frantar and Alistarh (2023) is not explicitly designed for output-balanced pruning granularity. When converting into an N:M sparsity pattern, SparseGPT forces every M consecutive weight in each row to have exactly N zeros. In such cases, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2024) unanimously remove the same amount of weights for each output. Such methods often emphasize the significance of individual components within the weight matrix and overlook inter-dependencies to other layers in the network. For MLP input projections pruning, an equal amount of weights corresponding to each intermediate neuron are removed. However, for output projection based on SparseGPT and Wanda pruning metrics, the weights connecting to intermediate neurons with larger activations are more likely to be reserved leading to a structural mismatch.
4 Dependency-aware Semi-structured Sparsity
In this section, we introduce Dependency-aware Semi-structured Sparsity (DaSS) for pruning MLP, which incorporates structural dependency into weight magnitude-based unstructured pruning method. An overview of DaSS and its comparison with the existing pruning method is shown in Figure 1.
Here, we denote the transposed weight matrices $\mathbf{W_{1}},\mathbf{W_{2}},\mathbf{W_{3}}$ in Eqs. (2) and (3) as $\mathbf{W^{(1)}}$, $\mathbf{W^{(2)}}$ and $\mathbf{W^{(3)}}$ , respectively. For $\mathbf{W^{(1)}}$ and $\mathbf{W^{(2)}}$, the shapes are $(d_{\text{int}},d_{\text{hidden}})$. For $\mathbf{W^{(3)}}$, the shape is $(d_{\text{hidden}},d_{\text{int}})$.
Structure Dependency in MLP.
In dependency-based structured pruning Ma et al. (2023); Fang et al. (2023), the initial step is dedicated to recognizing groups of interconnected structures within the model. In terms of GLU-based MLP module pruning, there are three projection matrices, all weights connecting to an identical intermediate neuron are collectively classified into the same dependency group. When pruning weights in $\mathbf{W^{(1)}_{i,:}}$, it is essential to also prune all the weights in $\mathbf{W^{(2)}_{i,:}}$ and $\mathbf{W^{(3)}_{:,i}}$ to maintain the structural consistency of the pruned neural network. In DaSS pruning, instead of aggressively pruning all the weights in less important dependency groups, we incorporate such structural coordination into unstructured pruning. If all weights in $\mathbf{W^{(3)}_{:,i}}$ are emphasized importance through additional importance indicator, such as augmenting corresponding input activations in Eq.(1), then the weights in $\mathbf{W^{(1)}_{i,:}}$ and $\mathbf{W^{(2)}_{i,:}}$ should also be emphasized similar importance.
Incorporating Dependency into Weight Importance.
In dependency-based pruning, we assess the importance of each weight and then aggregate the importance scores within the same group as the group importance score. Consequently, each weight within the same dependency group shares a consistent importance score, ensuring their simultaneous retention or elimination. The importance score of each weight is equal to the group importance score. In DaSS pruning, we take both group importance and weight magnitude into consideration to evaluate the importance of each weight. LLM-pruner Ma et al. (2023) evaluates the group importance using gradient-based methods. However, computing gradient for LLMs will introduce a significant amount of memory cost and it is less practical for larger models. The existing unstructured pruning methods Sun et al. (2024); Frantar and Alistarh (2023) are more efficient than gradient-based methods. By prioritizing the weights linked to outliers in intermediate activations, Wanda outperforms magnitude pruning in Down-Proj pruning. To minimize the impact on these critical outliers, it would be beneficial to also give greater significance to the weights that lead to outlier generation in both Gate-Proj and Up-Proj projections. Thus, we use the norm of intermediate activations $\|\mathbf{y}\|_{2}$ as the group importance indicator. The computation of activation norms is straightforward without introducing more computation and memory overhead than computing gradient. To assign larger importance scores on weights corresponding to more important groups, we evaluate weight importance based on the product of weight magnitude and corresponding group importance score. More specifically, for the Gate-Proj and Up-Proj , the importance score attributed to a specific weight $\mathbf{W}_{i,j}^{k}$ is determined as follows:
$\mathbf{I}_{i,j}^{k}=\left|\mathbf{W}_{i,j}^{k}\right|\cdot\|\mathbf{y}_{i}\|_{2}^{\alpha}$ (4)
where $k=1,2$. We introduce a hyper-parameter group importance strength $\alpha$. Our empirical results show that $\alpha=0.5$ is a well-balanced point for different models and datasets. For the Down-Proj projection pruning, we directly augment group importance into weight magnitude without additional hyper-parameter. For Down-Proj matrix, the importance score of weight $\mathbf{W}_{i,j}^{3}$ is determined as follows:
$\mathbf{I}_{i,j}^{3}=\left|\mathbf{W}_{i,j}^{3}\right|\cdot\|\mathbf{y}_{j}\|_{2}$ (5)
In Down-Proj pruning, the pruning metric is the same with Wanda Sun et al. (2024). By augmenting intermediate activations into all three weight importance matrices, DaSS inherently assigns greater emphasis to weights corresponding to intermediate activation outliers, thereby facilitating more nuanced structural coordination among the entire MLP module.
Pruning Granularity.
Pruning LLaMA models in finer granularity can improve the performance Sun et al. (2024). To incorporate intermediate activations into SwiGLU-based MLP pruning, each weight in the same comparison group should correspond to different intermediate activations. In Section 2.3, we show that input-balanced magnitude pruning can still achieve competitive results for Gate-Proj and Up-Proj pruning. In such pruning granularity, we can augment intermediate activations into Gate-Proj and Up-Proj pruning metric. In input-balanced pruning, we remove $s\%$ of the weights linked to each input for a predetermined sparsity ratio of $s\%$ based on weight importance scores. DaSS uses output-balanced sparsity for Down-Proj pruning, which is the same as Wanda.
Extension to N:M Sparsity.
The DaSS pruning design allows for easy adaptation to the N:M sparsity pattern. For Gate-Proj and Up-Proj the N: M sparsity pattern is formed on an input-balanced basis. This means that for weights connecting to each input neuron, out of every group of M consecutive weights, there are exactly N zeros included. For Down-Proj the N: M sparsity pattern is formed on an output-balanced basis.
Discussion.
In summary, our DaSS method offers multiple appealing aspects for pruning LLMs:
  1. 1.
    It retains the fundamental simplicity inherent in Wanda pruning method. Without weight updating, it still matches the performance of SparseGPT even at high sparsity as demonstrated in Section 4.4. This demonstrates the consistently effective and efficient capability of the DaSS method in identifying sparse neural networks.
  2. 2.
    Unlike SparseGPT and Wanda that use input + intermediate activations for MLP pruning, DaSS only uses intermediate activations. By using intermediate activations as group importance indicator, DaSS prunes MLP module in a more comprehensive view that captures the collective importance of all the weights connecting to each intermediate neuron.
  3. 3.
    DaSS effectively explores the balance between unstructured pruning’s flexibility and the structural coherence in dependency-based structured pruning.
PPL ($\downarrow$)
MLP Sparsity Method Gemma (GeGLU) ReluLLaMA (ReGLU) Mistral (SwiGLU) LLaMA2 (SwiGLU)
7B 7B 13B 7B 7B 13B 70B
Dense - 7.00 6.15 5.50 5.25 5.47 4.88 3.32
4:8 SparseGPT 10.86 8.44 7.45 7.36 7.33 6.29 4.66
Wanda 14.41 9.21 7.74 7.38 7.63 6.42 4.59
DaSS 10.86 8.19 7.16 7.06 7.26 6.16 4.41
2:4 SparseGPT 13.93 10.26 8.98 8.86 8.72 7.30 5.32
Wanda 32.50 12.68 9.87 9.24 9.55 7.68 5.35
DaSS 13.69 9.61 8.18 8.39 8.48 6.90 4.91
50% SparseGPT 9.64 7.22 6.39 6.20 6.38 5.60 4.06
Wanda 9.93 7.52 6.53 6.25 6.50 5.67 4.07
DaSS 9.11 7.24 6.40 6.15 6.44 5.59 4.00
Table 2: WikiText perplexity of pruned LLMs. The (*) in Size indicates a model that uses larger MLP than Vallina Transformer (e.g., grouped query attention).
5 Experiments
5.1 Settings
Models.
DaSS’s performance is evaluated over open LLMs using GLU variants. SwiGLU is the most widely used GLU-based MLP in the recent LLMs, including the LLaMA2 model family Touvron et al. (2023), which has models with parameters ranging between 7 billion and 70 billion, and also the Mistral-7B model Jiang et al. (2023). Among them, LLaMA2-70B and Mistral-7B use grouped-query attention Ainslie et al. (2023), the MLP module parameter accounts for around 80% of the total model parameters. There are only a few open LLMs that use other variants. For GeGLU, we use Gemma-7B. It is worth noting that the MLP intermediate dimension of Gemma-7B is 8×8$\times$8 × model dimension, making the MLP module much larger. For ReGLU, we use ReluLLaMA Team (2023), which is fine-tuned using ReGLU variant Shazeer (2020); Mirzadeh et al. (2023) based on LLaMA2 with small accuracy loss. The model configuration details are in Appendix A.1. We access the public checkpoints of the involved models provided by HuggingFace Transformers Wolf et al. (2019).
Baseline Approaches.
We compare the performance with two LLM-specific one-shot pruning approaches, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2024). We don’t consider structured pruning methods like LLM-Pruner Ma et al. (2023) and Sheared LLaMA Xia et al. (2023b), as they typically require re-training to recover accuracy, and are less practical for large models like LLaMA2-70B. Those baseline methods utilize uniform layerwise sparsity that can be easily converted into hardware-friendly N:M sparsity pattern. We used the same calibration data set as SparseGPT and Wanda in their model pruning processes, consisting of 128 sequences of 2048 tokens each, randomly selected from the first shard of the C4 dataset Raffel et al. (2020).
Evaluation.
To comprehensively evaluate the efficacy of our proposed method, two unique metrics are utilized to evaluate the performance of the pruned models: (1) perplexity (PPL) of language modeling (2) zero-shot accuracy on 5 commonsense reasoning tasks. Perplexity has been regarded as a consistent and reliable metric for measuring compressed models Dettmers and Zettlemoyer (2023); Frantar and Alistarh (2023), downstream tasks tend to be noisy, but more interpretable. For perplexity evaluation, we use the validation dataset of WikiText2 Merity et al. (2017). For zero-shot commonsense reasoning tasks, we choose five widely used tasks for accuracy evaluation: ARC (Easy and Challenge) Clark et al. (2018), HellaSwag Zellers et al. (2019), PiQA Bisk et al. (2020), and WinoGrande Sakaguchi et al. (2021), implemented in the Lm-Evaluation-Harness Gao et al. (2021). We evaluate the perplexity of all the aforementioned models. To fully demonstrate the task-wise performance in different sparsity patterns, we report the downstream task performance of the largest LLaMA2-70B model.
Sparsity.
In the less interpretable perplexity evaluation, we only prune the MLP layers. In the task-wise evaluation of LLaMA2-70B model, we prune both attention and MLP modules in accordance with previous works to better understand the performance gap between the pruned and original models. For attention module pruning, we use Wanda method, which is more efficient than SparseGPT. In LlaMA2-70B, as the MLP module accounts for more than 80% total parameters, we observe the choice of attention pruning has almost no impact on the final performance. We apply a uniform sparsity ratio across all the pruned layers and evaluate three sparsity types: unstructured sparsity, and semi-structured sparsities of 4:8 and 2:4.
5.2 Language Modeling
This evaluation involved a detailed examination of all the aforementioned LLMs. In perplexity evaluation, as shown in Table 2, our method consistently achieves better performance than SparseGPT and Wanda in more constrained and practical N:M sparsity pattern. As indicated by Sun et al. (2024), where weight updates can improve the performance in N:M sparsity pattern, our method shows superior performance even without computationally expensive weight updates. For Mistral-7B, Gamma-7B and LLaMA2-70B models with larger MLP layers, our method also outperforms SparseGPT in unstructured sparsity. Impressively, for the Gemma-7B with much larger MLP layers, DaSS outperforms Wanda in N:M sparsity significantly. DaSS shows consistent effectiveness across SwiGLU, GeGLU, and ReGLU, proving the generalizability of the DaSS method on GLU variants.
5.3 Downstream Tasks
While perplexity has been a popular metric in earlier LLM compression works Frantar and Alistarh (2023); Dettmers and Zettlemoyer (2023), its result is less interpretable. Apart from assessing perplexity, we comprehensively evaluate the performance of pruned LLaMA2-70B models in the widely used commonsense reasoning tasks.
In Table 3, we present the performance of different sparse LLaMA2-70B models on downstream tasks with prompting. The results show that our method outperforms SparseGPT and Wanda in most tasks at semi-structured N:M sparsity pattern. The only exception is that SparseGPT outperforms both Wanda and DaSS in Winogrande task. 50% unstructured SparseGPT pruned model even outperforms the dense model. Such results align with the prevailing LLM compression works Dettmers and Zettlemoyer (2023); Frantar and Alistarh (2023), where single task results can be noisy, the mean accuracy provides stable and reliable results. For unstructured sparsity, our method outperforms Wanda sharing the same complexity level.
Sparsity Method CommonSenseQA ($\uparrow$)
PIQA HellaSwag Winogrande ARC-e ARC-c Average
Dense - 82.15 66.05 77.98 82.55 54.35 72.62
4:8 SparseGPT 80.52 61.00 77.03 79.85 50.60 69.80
Wanda 80.47 61.85 75.45 80.10 50.00 69.57
DaSS 80.79 62.70 76.09 81.20 51.19 70.39
2:4 SparseGPT 79.00 59.00 76.64 78.50 47.87 68.20
Wanda 79.22 59.25 74.66 78.90 47.01 67.81
DaSS 79.70 60.00 74.82 79.65 49.15 68.66
50% SparseGPT 81.50 64.20 78.45 81.90 52.73 71.76
Wanda 81.01 64.30 77.35 80.95 52.05 71.13
DaSS 81.18 64.60 77.90 81.35 51.71 71.35
Table 3: Downstream tasks performance of LLaMA2-70B model in different sparsity pattern.
5.4 Performance Analysis
Sparsity Variation
MLP Sparsity 40% 50% 60% 70% 80%
SparseGPT 72.28 71.97 70.08 65.53 52.14
Wanda 72.35 71.49 69.47 62.75 40.76
DaSS 72.51 71.89 70.05 66.38 53.00
Table 4: LLaMA2-70B mean zero-shot tasks performance at different MLP sparsity ratios
Table 4 illustrates a comparison of the mean zero-shot task accuracy at varying levels of sparsity for the MLP module of the LLaMA2-70B model. It’s evident that DaSS pruning maintains competitive performance which closely matches that of SparseGPT across the entire range of sparsity ratios tested. Notably, DaSS achieves such performance without the need for weight updates, suggesting its effectiveness and efficiency in locating sparse neural networks. On the other hand, outlier-dependent Wanda pruning method shows a significant decline in accuracy as the sparsity ratio increases. This clear empirical evidence challenges the prevailing notion Dettmers et al. (2022); Lin et al. (2024) that weights corresponding to outliers are more important.
5.5 Running Time Analysis
Pruning speed
For DaSS and Wanda, the computational complexity is quantified as $O(d^{2})$, while SparseGPT exhibits a higher complexity of $O(d^{3})$. We recorded the overall time taken for pruning MLP layers, not including the forward pass process, in accordance with the approach described by Sun et al. (2024). We use a single A6000 48GB GPU to prune the 7B and 13B models, and use 8 A100 40GB GPUs to prune the larger 70B model.
Method Mistral LLaMA-2
7B 7B 13B 70B
SparseGPT 155 185 206 1092.76
Wanda 0.60 0.54 0.85 14.13
DaSS 0.81 0.81 1.02 20.70
Table 5: Pruning speed (in seconds) comparison.
As demonstrated in Table 5, the computational overhead incurred by DaSS is minimal, especially when compared to SparseGPT. Although the processing speed of DaSS is marginally slower than Wanda, this difference arises primarily from the implementation details of torch.sort(). In the implementation of torch.sort(), sorting along the last dimension is more efficient. Therefore, the observed delay in DaSS is attributed to the sorting operations rather than an inherent inefficiency in our method. Despite this, the substantial improvements in accuracy achieved by DaSS make the additional computational time justifiable and beneficial.
Sparsity Dense 50% 70%
Tokens per seconds 3.05 5.64 8.27
Speedup 1.0×1.0$\times$1.0 × 1.85×1.85$\times$1.85 × 2.71×2.71$\times$2.71 ×
Table 6: Decode inference performance for varying sparsity levels with LLaMA2-7B utilizing DeepSparse NeuralMagic (2021) inference engine.
Inference efficiency
We prune both attention and MLP module to test speed up. The speedup achieved by sparse model, as demonstrated in Table 6, highlights a significant reduction in end-to-end decode latency when utilizing LLaMA2-7B on the DeepSparse inference engine NeuralMagic (2021). The speed is tested on an Intel Xeon Platinum 8160 CPU with 24 cores. We highlight that the inference speedup is not exclusive to our pruning method but is a result of the inherent power of sparsity in accelerating computation.
6 Conclusion
Our observations reveal a diminishing importance of outliers in GLU-based MLP modules. Based on this insight, we propose the Dependency-aware Semi-structured Sparsity (DaSS) method. DaSS effectively balances the adaptability of unstructured pruning with the structural consistency of dependency-based structured pruning, utilizing MLP intermediate activation norms as the group importance indicator to develop a novel pruning metric. Empirical evaluations on the Mistral, Gemma and LLaMA2 model families show that DaSS surpasses state-of-the-art pruning methods such as SparseGPT and Wanda in achieving hardware-friendly N:M sparsity patterns. We hope our work inspires further research and development in the compression of GLU-based LLMs.
Limitations
Our research has the following limitations:
Merits beyond inference efficiency
Beyond improving inference efficiency, pruning techniques have demonstrated additional benefits for LLMs. Specifically, pruned models tend to exhibit fewer hallucinations compared to their unpruned counterparts Chrysostomou et al. (2024). Additionally, pruning has been shown to enhance resilience against jailbreak prompting Hasan et al. (2024) and to help in reducing toxicity and bias Xu and Hu (2022). However, the potential advantages of DaSS pruning beyond inference efficiency remain unclear and warrant further investigation.
Speed test in GPU
For LLM inference, the primary contributor to inference latency is the token generation process. While PyTorch’s current 2:4 sparsity implementation is optimized for accelerating matrix multiplication, it does not significantly enhance the performance of token generation. Recent work nm-vllm Shaw and Goin (2024) has shown approximately 1.5×1.5$\times$1.5 × token generation speedup for LlaMA2-7B at 50% unstructured sparsity. However, since this library is newly released and still under active development, it currently has many bugs, and we have not yet successfully implemented it. Kurtic et al. (2023) reports the speedup of several N:M sparsity patterns, its implementation hasn’t been released yet. The hardware and software support for sparsity is rapidly advancing, which will make sparse models more user-friendly in the future.
Application to other model architectures
Our method leverages a newly discovered property of GLU-based MLPs, making it unsuitable for older LLMs, such as OPT Zhang et al. (2022), which do not incorporate GLU variants. Additionally, our approach does not extend to the attention module, as the outlier activations still play an important role in the attention module.
Appendix A Appendix
A.1 Model Configurations
Table 11 is the configurations of the models used in the paper. We did not use LLaMA2-34B as it was not released. ReluLLaMA uses the same configuration as LLaMA2, with the only difference being the activation function. The links to the ReluLLaMA models are provided below:
A.2 Robustness to calibration samples.
Ashkboos et al. (2023) observes that the intermediate activations of SwiGLU-based MLP layers exhibit high variance, primarily caused by the Hadamard product of the preceding two outputs. This leads to diminished accuracy in 4-bit quantization. In Figure 2, we present how varying the number of sequences sampled for calibration affects the performance of pruning methods. Even though only using intermediate activations, our method demonstrates minimal sensitivity to changes in the number of calibration samples.
Figure 2: Robustness to calibration samples.
A.3 GeLU-based MLP Input projection
In the older MLP module, unlike GLU variants, there are only two projection matrices. We investigate magnitude pruning in different granularity for the MLP input projection matrix. We use the Falcon-7B model, which uses GeLU-based MLP. "Layer Magnitude" means we group weights in the entire weight matrix using magnitude pruning. In Table 7, we surprisingly find the input-balanced magnitude pruning achieves much worse results. This is in contrast to the results in GLU variants. The results show the huge significance of input activations for GeLU-based input projection pruning. Since our DaSS method needs to group weights per input, our method is less applicable to older MLP modules. However, GeLU-based MLP is rarely used in the current state-of-the-art LLMs.
Sparsity 50% 60% 70%
Wanda 7.17 8.72 18.95
Output-balanced Magnitude 8.46 15.27 498.77
Input-balanced Magnitude 1708 10587 13706
Layer Magnitude 21.82 1895 12896
Table 7: The Wikitext perplexity results of GeLU input projection at different sparsity ratio
A.4 Combination with Quantization
Method SaprseGPT Wanda DaSS
50% sparsity 6.38 6.50 6.44
w/ 4-bit AWQ 6.51 6.63 6.57
Table 8: The Wikitext perplexity results of LLaMA2-7B with 50% sparsity and 4 bit quantization
Pruning and quantization, traditionally seen as distinct model compression techniques, are not mutually exclusive and can be effectively integrated to enhance overall efficiency Han et al. (2015a); Frantar and Alistarh (2023); Agarwalla et al. (2024). Here we test the combination of DaSS with 4-bit AWQ Lin et al. (2024) for compressing LlaMA2-7B. As shown in Table 8, the performance of all the pruned models just drops slightly after 4-bit quantization. DaSS still outperforms wanda after 4-bit quantization.
A.5 MMLU Results
Concurrent work (Gromov et al., 2024) indicates the unusual behavior of LLMs in performing Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021) tasks. We use Chain-of-Thought Hub Fu et al. (2023) which is based on the official implementation of MMLU Hendrycks et al. (2021). The MMLU encompasses 57 tasks, spanning from STEM, Humanities, Social Sciences, among others and we report the mean accuracy of 57 tasks. We test the performance of pruned LlaMA2-70B model in MMLU task.
Sparsity Methods MMLU (5 shot)
Dense - 69.10
4:8 SparseGPT 60.24
Wanda 59.69
DaSS 60.86
DaSS+skip 1/4 65.82
2:4 SparseGPT 56.99
Wanda 56.41
DaSS 57.53
DaSS+skip 1/4 64.36
50% SparseGPT 64.52
Wanda 62.72
DaSS 63.27
DaSS+skip 1/4 66.81
Table 9: MMLU scores for different sparsity methods and levels
From Table 9, we observe that the improvement of DaSS over Wanda becomes more pronounced in the challenging MMLU task, where it achieves an increase in accuracy of 1.17 and 1.12 for the 4:8 and 2:4 sparsity patterns, respectively. In accordance with the observations with Jaiswal et al. (2023), we see a considerable performance degradation in knowledge-intensive MMLU task, in which only unstructured 50% sparsity models outperform the dense LLaMA2-34B model. As GPU plays a more important role in larger model inference, it is essential to improve the performance of pruned models in hardware-friendly N:M sparsity. Uniform reduction of the overall sparsity level is not feasible for N:M sparsity, Frantar and Alistarh (2023) suggests a specific subset of layers can be chosen for full N:M sparsification. Here we skip pruning 1/4 consecutive layers (20 layers) and result in a final 37.5% sparsity ratio. We use the first 10 tasks of MMLU to study pruning sensitivity. We divide the model into 4 consecutive parts and study skipping pruning each part. As shown in Table 10, the earlier layers are more sensitive than the later ones in knowledge-intensitive tasks, which is contradictory to the findings in Frantar and Alistarh (2023) using perplexity, and align with the findings of Gromov et al. (2024).
Start Index 0 10 20 40 60 Dense
Acc (%) 60.12 61.72 59.75 56.54 56.83 64.86
Table 10: MMLU subset accuracy after skipping pruning 20 layers at various start indices in 4:8 sparsity.
Model Param Layers Hidden Intermediate Query Heads KV Head
LlaMa2-7B 7B 32 4096 11008 32 32
LlaMa2-13B 13B 40 5120 13824 40 40
LlaMa2-70B 70B 80 8192 28672 64 8
Mistral-7B 7B 32 4096 14336 32 8
Gemma-7B 7B 28 3072 24576 16 16
Table 11: Model configurations of Llama2, Gemma and Mistral models.
We continue to search for the better skipping layers with the start layer index range in [0,20]. We found that starting skipping from layer 10 can achieve the best performance in the subset. Then we test its results in the full MMLU tasks. As shown in Table 3, skipping sensitive 1/4 layers can significantly improve the performance of pruned models, especially for N:M sparsity. We can achieve sparse models that perform better than LLaMA2-34B. Although partial N:M sparsity models have more parameters than smaller dense models, training many smaller dense models like LLaMA2-34B is still computationally expensive. Efficient post-training pruning enables us to easily adjust the accuracy-efficiency trade-off in real-world applications.
Concurrent work Gromov et al. (2024) propose an effective method to remove half of the layers in LlaMA2-70B with small accuracy loss for MMLU tasks, suggesting the unusual behavior of LLMs in performing MMLU tasks. Their work is focused more on task-specific pruning, while our work emphasizes preserving the general ability of LLMs to perform various tasks. Our weight pruning method is fundamentally orthogonal to layer pruning in principle.